OpenStudy (anonymous):

20g^2=16g Can you show your work?

OpenStudy (anonymous):

divide by 4 and get \[5g^2=4g\] subtract to get \[5g^2-4g=0\] and then factor to solve \[g(5g-4)=0\]

OpenStudy (anonymous):

Thanks so much! :)

OpenStudy (anonymous):

yw